Given that Point A (3,4) lies on the graph of g(x), and Point A' (3,8) lies on the graph of h(x), where h(x)=(f o g) (x), the corresponding point that lies on the graph of f(x) must be ____
The teacher answer is (4,8) BUT WHY???

Answer by Theo(13342)


the point (3,4) means that g(x) = 4 when x = 3.

the point (3,8) means that h(x) = 8 when x = 3.

you are given that h(x) = fog(x).

by definition fog(x) = f(g(x)).

when x = 3, g(x) = 4, which means that f(g(x)) = f(4) = 8.

f(4) = 8 results in the point (4,8).